a father is 3 times as old as his son. 15 years ago, the father was 9 times as old as his son was then. Find thier present ages.
Answer by scott8148(6628)

"a father is 3 times as old as his son" ___ f = 3s

"15 years ago, the father was 9 times as old as his son was then" ___ f-15 = 9(s-15)

substituting ___ 3s - 15 = 9(s-15)

distributing ___ 3s - 15 = 9s - 135

adding 135-3s ___ 120 = 6s

dividing by 6 ___ 20 = s

substituting ___ f = 3(20) = 60